# On Sobolev and Besov spaces with hybrid regularity

### v1

@@ 5.1. Change of Bases
First, we need to find a way to express the isotropic basis functions in terms of tensor-product functions. For simplicity, we only treat the case $n=2$ explicitly again, but we emphasise that the main results, Theorems 5.4 and 5.6, carry over to the $n$-dimensional case as well, if $2s$ is replaced by $sn$.
@@
$|u|_{\mathfrak{B}^{q,s,\tau}_{\tau}(\square)}^{\tau}$ $=|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{q,s,\tau}_{\tau}}^{\tau}=\sum_{\mathbf{j}\geq\mathbf{j}_{0}}2^{\tau q|\mathbf{j}|_{\infty}+\tau s|\mathbf{j}|_{1}}\|\mathbf{u}^{(\tau)}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}$
@@
$|u|_{B^{\alpha,\tau}_{\tau}(\square)}^{\tau}$ $=|\mathbf{v}^{(\tau)}|_{\mathbf{b}^{\alpha,\tau}_{\tau}}^{\tau}=\sum_{m=j_{0}}^{\infty}2^{\tau\alpha m}\|\mathbf{v}^{(\tau)}\|_{\ell^{\tau}(\hexagon_{m})}^{\tau}$
@@
$=\sum_{m=j_{0}}^{\infty}2^{\tau m(\alpha-2s)}\left(\big{\|}(\mathbf{T}_{m-1}\otimes\mathbf{I})\mathbf{u}\big{|}_{\bigcup_{j<m}\nabla_{j}\times\nabla_{m}}\big{\|}_{\ell^{\tau}(\hexagon_{m})^{(0,1)}}^{\tau}\right.$
$\qquad+\left.\big{\|}(\mathbf{I}\otimes\mathbf{T}_{m-1})\mathbf{u}\big{|}_{\bigcup_{j<m}\nabla_{m}\times\nabla_{j}}\big{\|}_{\ell^{\tau}(\hexagon_{m})^{(1,0)}}^{\tau}+\|\mathbf{u}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{(m,m)})}^{\tau}\right).$ (26)
@@ 5.2. Comparison of Approximation Spaces
The goal of this section is to compare the classical approximation spaces $B^{q+2s,\tau}_{\tau}(\square)$ with the approximation spaces for tensor-product wavelets $\mathfrak{B}^{q,s,\tau}_{\tau}(\square)$. We will see that there is a range of regularity spaces whose elements can be approximated by tensor-product wavelets but not by isotropic wavelets.
@@
$|u|_{\mathfrak{B}^{s,0,\tau}_{\tau}(\square)}\lesssim\|u\|_{B^{s+\varepsilon,\tau}_{\tau}(\square)}.$
@@
We remark that we need to consider Besov spaces with logarithmic correction, cf. [1]. Due to [1, Theorem 2.6] (after rescaling the coefficients such that they correspond to wavelets normalised in $L^{\tau}(\square)$), there holds $u=\Psi\mathbf{u}\in B^{s,\tau}_{\tau,|\log|^{\beta}}$ if
$\overline{\sigma}^{s}(\mathbf{u})\mathrel{\mathrel{\mathop{:}}=}\left[\sum_{\mathbf{j}\geq\mathbf{j}_{0}}|\mathbf{j}|_{\infty}^{\overline{\varrho}-\beta\tau}2^{\tau s|\mathbf{j}|_{\infty}}\|\mathbf{u}^{(\tau)}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}\right]^{\frac{1}{\tau}}<\infty.$
On the other hand, if $u=\Psi\mathbf{u}\in B^{s,\tau}_{\tau,|\log|^{\beta}}$, then
$\underline{\sigma}^{s}(\mathbf{u})\mathrel{\mathrel{\mathop{:}}=}\left[\sum_{\mathbf{j}\geq\mathbf{j}_{0}}|\mathbf{j}|_{\infty}^{-\underline{\varrho}-\beta\tau-1}2^{\tau s|\mathbf{j}|_{\infty}}|\mathbf{u}^{(\tau)}|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}\right]^{\frac{1}{\tau}}<\infty.$
Herein,
$\beta=\left|\frac{1}{\tau}-1\right|,\qquad\overline{\varrho}=\begin{cases}1-\tau,&s\geq\frac{1}{2},\\ 0,&s<\frac{1}{2},\end{cases}\qquad\underline{\varrho}=\begin{cases}1,&s>\frac{1}{2},\\ \max\{2-\tau,0\},&s\leq\frac{1}{2}.\end{cases}$
Moreover, as $\beta\geq 0$, there holds $B^{s+\varepsilon,\tau}_{\tau}(\square)\hookrightarrow B^{s+\varepsilon,\tau}_{\tau,|\log|^{\beta}}(\square)\hookrightarrow B^{s,\tau}_{\tau}(\square)$. Hence, if we can show that $|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{s,0,\tau}_{\tau}}\lesssim\underline{\sigma}^{s+\varepsilon}(\mathbf{u})$, we have the embeddings
$B^{s+\varepsilon,\tau}_{\tau}(\square)\hookrightarrow B^{s+\varepsilon,\tau}_{\tau,|\log|^{\beta}}(\square)\hookrightarrow\mathfrak{B}^{s,0,\tau}_{\tau}(\square),$
which is what we want to show.
To this end, let $\mathbf{u}^{(\tau)}\in\mathfrak{b}^{s,0,\tau}_{\tau}$. By standard estimates, there holds
$|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{s,0,\tau}_{\tau}}^{\tau}$ $=\sum_{\mathbf{j}\geq\mathbf{j}_{0}}2^{\tau s|\mathbf{j}|_{\infty}}\|\mathbf{u}^{(\tau)}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}\lesssim\sum_{\mathbf{j}\geq\mathbf{j}_{0}}|\mathbf{j}|_{\infty}^{-\underline{\varrho}-\beta\tau-1}2^{\tau(s+\varepsilon)|\mathbf{j}|_{\infty}}\|\mathbf{u}^{(\tau)}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}$
$=\underline{\sigma}^{s+\varepsilon}(\mathbf{u})^{\tau},$
by which this theorem is proven. ∎
Remark 5.2.
By the same arguments, since $\overline{\varrho}-\beta\tau<0$, we also see that we have
$\|u\|_{B^{s,\tau}_{\tau,|\log|^{\beta}}(\square)}\lesssim|u|_{\mathfrak{B}^{s,0,\tau}_{\tau}(\square)}.$
Corollary 5.3.
$|u|_{\mathfrak{B}^{q,s,\tau}_{\tau}(\square)}\lesssim|u|_{\mathfrak{B}^{q+2s,0,\tau}_{\tau}(\square)}\lesssim\|u\|_{B^{q+2s+\varepsilon,\tau}_{\tau}(\square)}.$
On the other hand, there also holds
$\|u\|_{B^{q+s-\varepsilon,\tau}_{\tau}(\square)}\lesssim|u|_{\mathfrak{B}^{q+s,0,\tau}_{\tau}(\square)}\lesssim|u|_{\mathfrak{B}^{q,s,\tau}_{\tau}(\square)}.$
Proof.
The first embedding is due to Theorem 5.1, whereas the second embedding follows trivially from the definition of the seminorm $|\cdot|_{\mathfrak{b}^{q,s,\tau}_{\tau}}$ in (18) and the estimate $|\mathbf{j}|_{\infty}\leq|\mathbf{j}|_{1}\leq 2|\mathbf{j}|_{\infty}$. ∎
This corollary immediately implies that there are a lot of functions which satisfy $u\in\mathfrak{A}^{s}\big{(}H^{q}(\square)\big{)}$ but $u\notin A^{s}\big{(}H^{q}(\square)\big{)}$, where $\mathfrak{A}^{s}$ and $A^{s}$ denote the approximation spaces with respect to the tensor-product and the isotropic wavelet bases, respectively. In particular, all functions which admit slightly more regularity than minimally required to be in $A^{s}\big{(}H^{q}(\square)\big{)}$ are also in $\mathfrak{A}^{s}\big{(}H^{q}(\square)\big{)}$. Nevertheless, there might also exist functions $u\in A^{s}\big{(}H^{q}(\square)\big{)}$, with $u\notin\mathfrak{A}^{s}\big{(}H^{q}(\square)\big{)}$, i.e., functions which can be approximated better by isotropic wavelets.
In the case of wavelets with compactly supported duals, however, as the following theorem shows, this is not possible.
Theorem 5.4.
For $u=\Psi\mathbf{u}=\Theta\mathbf{v}$, where $\Psi$ and $\Theta$ are the tensor-product and isotropic wavelet bases defined in Section 2.2, there holds
$|\mathbf{v}^{(\tau)}|_{\mathbf{b}^{q+s,\tau}_{\tau}}\lesssim|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{q,s,\tau}_{\tau}}\lesssim|\mathbf{v}^{(\tau)}|_{\mathbf{b}^{q+2s,\tau}_{\tau}},$
with $\frac{1}{\tau}=s+\frac{1}{2}$. In particular, we have $|u|_{B^{q+s,\tau}_{\tau}(\square)}\lesssim|u|_{\mathfrak{B}^{q,s,\tau}_{\tau}(\square)}\lesssim|u|_{B^{q+2s,\tau}_{\tau}(\square)}$.
Proof.
Let $u=\Psi\mathbf{u}=\Theta\mathbf{v}\in B^{q+2s,\tau}_{\tau}(\square)$. We first want to apply the suitable coordinate transforms to derive the estimate
$|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{q,s,\tau}_{\tau}}\lesssim|\mathbf{v}^{(\tau)}|_{\mathbf{b}^{q+2s,\tau}_{\tau}}=\left[\sum_{m=0}^{\infty}2^{\tau m(q+2s)}\sum_{\mu\in\hexagon_{m}}|v_{\mu}^{(\tau)}|^{\tau}\right]^{\frac{1}{\tau}}.$
@@
$|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{q,s,\tau}_{\tau}}^{\tau}$ $=\sum_{m=j_{0}}^{\infty}2^{\tau mq}\left(\|(\tilde{\mathbf{T}}_{m-1}^{\intercal}\otimes\mathbf{I})\mathbf{v}|_{\hexagon_{m}^{(0,1)}}\|_{\ell^{\tau}(\bigcup_{j<m}\nabla_{j}\times\nabla_{m})}^{\tau}\right.$
$\qquad+\left.\|(\mathbf{I}\otimes\tilde{\mathbf{T}}_{m-1}^{\intercal})\mathbf{v}|_{\hexagon_{m}^{(1,0)}}\|_{\ell^{\tau}(\bigcup_{j<m}\nabla_{m}\times\nabla_{j})}^{\tau}\right)+2^{\tau mq}\|\mathbf{v}\|_{\ell^{\tau}(\hexagon_{m}^{(1,1)})}^{\tau}.$
@@
which can be shown to be uniformly bounded using the arguments of the proof of Lemma 2.3. If $1\leq\tau\leq 2$, and $\frac{1}{\tau}=\frac{1-\theta}{1}+\frac{\theta}{2}$, interpolation and Lemma 2.5 yield
@@
$|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{q,s,\tau}_{\tau}}^{\tau}$ $\lesssim\sum_{m=j_{0}}^{\infty}2^{\tau mq}\|\mathbf{v}\|_{\ell^{\tau}(\hexagon_{m})}^{\tau}=\sum_{m=j_{0}}^{\infty}2^{\tau m(q+2s)}\|\mathbf{v}^{(\tau)}\|_{\ell^{\tau}(\hexagon_{m})}^{\tau},$
@@
$|\mathbf{v}^{(\tau)}|_{\mathbf{b}^{q+s,\tau}_{\tau}}$ $=\sum_{m=j_{0}}^{\infty}2^{\tau m(q-s)}\left(\big{\|}(\mathbf{T}_{m-1}\otimes\mathbf{I})\mathbf{u}\big{|}_{\bigcup_{j<m}\nabla_{j}\times\nabla_{m}}\big{\|}_{\ell^{\tau}(\hexagon_{m})^{(0,1)}}^{\tau}\right.$
$\qquad+\left.\big{\|}(\mathbf{I}\otimes\mathbf{T}_{m-1})\mathbf{u}\big{|}_{\bigcup_{j<m}\nabla_{m}\times\nabla_{j}}\big{\|}_{\ell^{\tau}(\hexagon_{m})^{(1,0)}}^{\tau}+\|\mathbf{u}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{(m,m)})}^{\tau}\right).$ (27)
Also here, by using the arguments of the Lemmata 2.1, 2.5, and 2.3, we see that
@@
if $0<\tau\leq 1$. If $1\leq\tau\leq 2$ and $\frac{1}{\tau}=\frac{1-\theta}{1}+\frac{\theta}{2}$, meaning that $1-\theta=2s$, we may interpolate again to deduce that
@@
$|\mathbf{v}^{(\tau)}|_{\mathbf{b}^{q+s,\tau}_{\tau}}\lesssim\sum_{m=j_{0}}^{\infty}2^{\tau mq}\sum_{|\mathbf{j}|_{\infty}=m}\|\mathbf{u}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}=\sum_{\mathbf{j}\geq\mathbf{j}_{0}}2^{\tau q|\mathbf{j}|_{\infty}+\tau s|\mathbf{j}|_{1}}\|\mathbf{u}^{(\tau)}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}.$
@@
The most important step in the proof of Theorem 5.4 is the estimate of the transformation matrices. Although this was only carried out for the two-dimensional case, the same can be concluded in $n$ dimensions. Indeed, since if Lemma $\|(\tilde{\mathbf{T}}_{m}^{\odot\tau})^{\intercal}\|_{1}$ is uniformly bounded, Lemma 2.5 can be applied recursively.
For the opposite estimate, we need to be slightly more careful. Indeed, the matrices arising from the transforms may be given by $(n-1)$ tensor factors $\mathbf{T}_{m-1}$ and only one identity, resulting in a norm of $2^{\tau m(n-1)s}$. However, by rescaling the coefficients from $L^{\tau}(\square)$ to $L^{2}(\square)$, the weight in (27) is now given by $2^{\tau m(q-(n-1)s)}$, which multiplies to $2^{\tau mq}$, as desired.
With this remark, we can immediately deduce the following result.
Theorem 5.6.
Let $q,s>0$. Then, for $\frac{1}{\tau}=s+\frac{1}{2}$, the embeddings $B^{q+sn,\tau}_{\tau}(\square)\hookrightarrow\mathfrak{B}^{q,s,\tau}_{\tau}(\square)\hookrightarrow B^{q+s,\tau}_{\tau}(\square)$ are continuous.
Proof.
Since $q>0$, there holds with Theorem 5.4, that
$\|u\|_{\mathfrak{B}^{q,s,\tau}_{\tau}(\square)}=\|u\|_{L^{p}(\square)}+|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{q,s,\tau}_{\tau}}\lesssim\|u\|_{L^{p}(\square)}+|\mathbf{v}^{(\tau)}|_{\mathbf{b}^{q+sn,\tau}_{\tau}}=\|u\|_{B^{q+sn,\tau}_{\tau}(\square)}.$
Similarly,
$\|u\|_{B^{q+s,\tau}_{\tau}(\square)}=\|u\|_{L^{p}(\square)}+|\mathbf{v}^{(\tau)}|_{\mathbf{b}^{q+s,\tau}_{\tau}}\lesssim\|u\|_{L^{p}(\square)}+|\mathbf{u}^{(\tau)}|_{\mathfrak{b}^{q,s,\tau}_{\tau}}=\|u\|_{\mathfrak{B}^{q,s,\tau}_{\tau}(\square)},$
which implies the claim. ∎

### v2

@@ 5.1. Change of Bases
First, we need to find a way to express the isotropic basis functions in terms of tensor-product functions. For simplicity, we only treat the case $n=2$ explicitly again, but we emphasise that the main result, Theorem 5.3, carries over to the $n$-dimensional case as well if $2s$ is replaced by $sn$.
@@
$\|u\|_{\mathfrak{B}^{q,s,\tau}_{\tau}(\square)}^{\tau}$ $=\|\mathbf{u}^{(\tau)}\|_{\mathfrak{b}^{q,s,\tau}_{\tau}}^{\tau}=\sum_{\mathbf{j}\geq\mathbf{j}_{0}}2^{\tau q|\mathbf{j}|_{\infty}+\tau s|\mathbf{j}|_{1}}\|\mathbf{u}^{(\tau)}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}$
@@
$\|u\|_{B^{\alpha,\tau}_{\tau}(\square)}^{\tau}$ $=\|\mathbf{v}^{(\tau)}\|_{\mathbf{b}^{\alpha,\tau}_{\tau}}^{\tau}=\sum_{m=j_{0}}^{\infty}2^{\tau\alpha m}\|\mathbf{v}^{(\tau)}\|_{\ell^{\tau}(\hexagon_{m})}^{\tau}$
@@
$=\sum_{m=j_{0}}^{\infty}2^{\tau m(\alpha-2s)}\left(\big\|(\mathbf{T}_{m-1}\otimes\mathbf{I})\mathbf{u}\big|_{\bigcup_{j<m}\nabla_{j}\times\nabla_{m}}\big\|_{\ell^{\tau}(\hexagon_{m}^{(0,1)})}^{\tau}\right$.
$\qquad+\left.\big\|(\mathbf{I}\otimes\mathbf{T}_{m-1})\mathbf{u}\big|_{\bigcup_{j<m}\nabla_{m}\times\nabla_{j}}\big\|_{\ell^{\tau}(\hexagon_{m}^{(1,0)})}^{\tau}+\|\mathbf{u}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{(m,m)})}^{\tau}\right).$ (28)
@@ 5.2. Comparison of Approximation Spaces
The goal of this section is to compare the classical approximation spaces $B^{q+sn,\tau}_{\tau}(\square)$ with the approximation spaces for tensor-product wavelets $\mathfrak{B}^{q,s,\tau}_{\tau}(\square)$. We will see that there is a range of regularity spaces whose elements can be approximated by tensor-product wavelets but not by isotropic wavelets.
@@
Proposition 5.1.
$\|u\|_{B^{q+s-\varepsilon,\tau}_{\tau}(\square)}\lesssim\|u\|_{\mathfrak{B}^{q,s,\tau}_{\tau}(\square)}\lesssim\|u\|_{B^{q+2s+\varepsilon,\tau}_{\tau}(\square)}.$
Remark 5.2.
Proposition 5.1 can be improved by replacing the $\varepsilon$ by a logarithmic correction weight. In order to do so, Besov spaces with logarithmic correction need to be introduced, cf. [1]. Moreover, note that the results [1] in a two-dimensional setting.
Proposition 5.1 immediately implies that there are a lot of functions which satisfy $u\in\mathfrak{A}^{s}\big(H^{q}(\square)\big)$ but $u\notin A^{s}\big(H^{q}(\square)\big)$, where $\mathfrak{A}^{s}$ and $A^{s}$ denote the approximation spaces with respect to the tensor-product and the isotropic wavelet bases, respectively. In particular, all functions which admit slightly more regularity than minimally required to be in $A^{s}\big(H^{q}(\square)\big)$ are also in $\mathfrak{A}^{s}\big(H^{q}(\square)\big)$. In the next theorem, we will however show that no additional regularity is required.
@@
Suppose that the primal and dual scaling functions and wavelets satisfy the asymptotic decay estimate (11) for some $\alpha>1$. Moreover, let $s<\alpha-\frac{1}{2}$ and $q+s,q+sn>\max\{0,\,n(s-\frac{1}{2})\}$. Then, for $\frac{1}{\tau}=s+\frac{1}{2}$, the embeddings
$B^{q+sn,\tau}_{\tau}(\square)\hookrightarrow\mathfrak{B}^{q,s,\tau}_{\tau}(\square)\hookrightarrow B^{q+s,\tau}_{\tau}(\square)$
are continuous.
@@
Since $q+s,q+sn>\max\{0,\,n(\frac{1}{\tau}-1)\}$, both $B^{q+s,\tau}_{\tau}(\square)$ and $B^{q+sn,\tau}_{\tau}(\square)$ can be characterised by a sufficiently regular wavelet basis, cf. [4]. Therefore, it suffices to show that, for $u=\Psi\mathbf{u}=\Theta\mathbf{v}$,
$\|\mathbf{v}^{(\tau)}\|_{\mathbf{b}^{q+s,\tau}_{\tau}}\lesssim\|\mathbf{u}^{(\tau)}\|_{\mathfrak{b}^{q,s,\tau}_{\tau}}\lesssim\|\mathbf{v}^{(\tau)}\|_{\mathbf{b}^{q+sn,\tau}_{\tau}}.$
Let us first set $n=2$ and consider $u\in B^{q+2s,\tau}_{\tau}(\square)$. We want to apply a suitable coordinate transforms to derive the estimate
$\|\mathbf{u}^{(\tau)}\|_{\mathfrak{b}^{q,s,\tau}_{\tau}}\lesssim\|\mathbf{v}^{(\tau)}\|_{\mathbf{b}^{q+2s,\tau}_{\tau}}=\left[\sum_{m=0}^{\infty}2^{\tau m(q+2s)}\sum_{\mu\in\hexagon_{m}}\big|v_{\mu}^{(\tau)}\big|^{\tau}\right]^{\frac{1}{\tau}}.$
@@
$\|\mathbf{u}^{(\tau)}\|_{\mathfrak{b}^{q,s,\tau}_{\tau}}^{\tau}$ $=\sum_{m=j_{0}}^{\infty}2^{\tau mq}\left(\big\|(\tilde{\mathbf{T}}_{m-1}^{\intercal}\otimes\mathbf{I})\mathbf{v}|_{\hexagon_{m}^{(0,1)}}\big\|_{\ell^{\tau}(\bigcup_{j<m}\nabla_{j}\times\nabla_{m})}^{\tau}\right$.
$\qquad+\left.\big\|(\mathbf{I}\otimes\tilde{\mathbf{T}}_{m-1}^{\intercal})\mathbf{v}|_{\hexagon_{m}^{(1,0)}}\big\|_{\ell^{\tau}(\bigcup_{j<m}\nabla_{m}\times\nabla_{j})}^{\tau}\right)+2^{\tau mq}\|\mathbf{v}\|_{\ell^{\tau}(\hexagon_{m}^{(1,1)})}^{\tau}.$
@@
which can be shown to be uniformly bounded using the arguments of the proof of Lemma 2.5, as $s<\alpha-\frac{1}{2}$ ensures that $\tau>\frac{1}{\alpha}$. If $1\leq\tau\leq 2$, and $\frac{1}{\tau}=\frac{1-\theta}{1}+\frac{\theta}{2}$, interpolation and Lemma 2.6 yield
@@
$\|\mathbf{u}^{(\tau)}\|_{\mathfrak{b}^{q,s,\tau}_{\tau}}^{\tau}$ $\lesssim\sum_{m=j_{0}}^{\infty}2^{\tau mq}\|\mathbf{v}\|_{\ell^{\tau}(\hexagon_{m})}^{\tau}=\sum_{m=j_{0}}^{\infty}2^{\tau m(q+2s)}\|\mathbf{v}^{(\tau)}\|_{\ell^{\tau}(\hexagon_{m})}^{\tau},$
@@
$\|\mathbf{v}^{(\tau)}\|_{\mathbf{b}^{q+s,\tau}_{\tau}}$ $=\sum_{m=j_{0}}^{\infty}2^{\tau m(q-s)}\left(\big\|(\mathbf{T}_{m-1}\otimes\mathbf{I})\mathbf{u}\big|_{\bigcup_{j<m}\nabla_{j}\times\nabla_{m}}\big\|_{\ell^{\tau}(\hexagon_{m}^{(0,1)})}^{\tau}\right$.
$\qquad+\left.\big\|(\mathbf{I}\otimes\mathbf{T}_{m-1})\mathbf{u}\big|_{\bigcup_{j<m}\nabla_{m}\times\nabla_{j}}\big\|_{\ell^{\tau}(\hexagon_{m}^{(1,0)})}^{\tau}+\|\mathbf{u}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{(m,m)})}^{\tau}\right).$ (29)
Also here, by using the arguments of the Lemmata 2.1, 2.6, and 2.5, we see that
@@
because $\frac{1}{\alpha}<\tau\leq 1$. If $1\leq\tau\leq 2$ and $\frac{1}{\tau}=\frac{1-\theta}{1}+\frac{\theta}{2}$, meaning that $1-\theta=2s$, we may interpolate again to deduce that
@@
$\|\mathbf{v}^{(\tau)}\|_{\mathbf{b}^{q+s,\tau}_{\tau}}\lesssim\sum_{m=j_{0}}^{\infty}2^{\tau mq}\sum_{|\mathbf{j}|_{\infty}=m}\|\mathbf{u}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}=\sum_{\mathbf{j}\geq\mathbf{j}_{0}}2^{\tau q|\mathbf{j}|_{\infty}+\tau s|\mathbf{j}|_{1}}\|\mathbf{u}^{(\tau)}\|_{\ell^{\tau}(\boldsymbol{\nabla}_{\mathbf{j}})}^{\tau}.$
In the case $n\geq 2$, we apply Lemma 2.6 recursively to matrices arising from an $n$-fold tensor products. Since $\|(\tilde{\mathbf{T}}_{m}^{\odot\tau})^{\intercal}\|_{1}$ is uniformly bounded, the right estimate follows.
For the left estimate, we need to bound tensor products of at most (n1)(n-1) tensor factors $\mathbf{T}_{m-1}$ and at least one identity, resulting in a norm asymptotically bounded by $2^{\tau m(n-1)s}$. On the other hand, the weight in (29) results from the rescaling of the coefficients from $L^{\tau}(\square)$ to $L^{2}(\square)$, which gives rise to a factor $2^{\tau m(q-(n-1)s)}$ in $n$ dimensions. Hence, this multiplies to $2^{\tau mq}$, as in the two-dimensional case. ∎
@@
The maximal approximation rate $s$ is limited by the decay behaviour of the primal and dual scaling functions. However, if compactly supported wavelets are used, $\alpha$, and therefore also $s$, can be chosen arbitrarily large, by which this limitation can be mitigated. The same holds true for wavelets that decay faster than any polynomial.